how do you solve this problem????
A square vegetable garden is to be tilled and then enclosed with a fence. If the fence costs $1 per foot and the cost of preparing the soil is $0.50 per ft2 , determine the size of the garden that can be enclosed for $120.
i have a answer but i can't really follow this question...:((
could you explain it in details? :))

Answer by mananth(16946)

length of side of garden x ft.
perimeter = 4x
cost of fencing = 1*4x= 4x
..
Area = x*x = x^2
preparing ground = 0.5* x*x= 0.5x^2
...
0.5x^2+4x=120
multiply by 10
5x^2+40x-1200=0
/5
x^2+8x-240=0
x^2+20x-12x-240=0
x(x+20)-12(x+20)=0
(x+20)(x-12)=0
x= 12 which is positive.
The garden is 12 by 12 ft.